# Determining Energy Efficiency Sweet Spots in Production LLM Inference

### v1

@@ Abstract.
Large Language Models (LLMs) inference is central in modern AI applications, making it critical to understand their energy footprint. Existing approaches typically estimate energy consumption through simple linear functions of input and output sequence lengths, yet our observations reveal clear Energy Efficiency regimes: peak efficiency occurs with short-to-moderate inputs and medium-length outputs, while efficiency drops sharply for long inputs or very short outputs, indicating a non-linear dependency. In this work, we propose an analytical model derived from the computational and memory-access complexity of the Transformer architecture, capable of accurately characterizing the efficiency curve as a function of input and output lengths. To assess its accuracy, we evaluate energy consumption using TensorRT-LLM on NVIDIA H100 GPUs across a diverse set of LLMs ranging from 1B to 9B parameters, including OPT, LLaMA, Gemma, Falcon, Qwen2, and Granite, tested over input and output lengths from 64 to 4096 tokens, achieving a mean MAPE of 1.79%. Our results show that aligning sequence lengths with these efficiency “Sweet Spots” can substantially reduce energy usage, supporting informed truncation, summarization, and adaptive generation strategies in production systems.
LLM Energy Consumption, Joules per Token, Inference costs
@@ 1. Introduction
Large Language Models (LLMs) have rapidly emerged as the foundation of modern Natural Language Processing (NLP), enabling applications ranging from reasoning and summarization to dialogue systems and generative AI. Their adoption at scale, however, has introduced a growing concern: the substantial energy consumption associated with inference, which is estimated to account for up to 90% of the model’s consumption (Jegham et al., 2025). Unlike training, which is an episodic process, inference occurs continuously in production settings, every time an LLM generates output for a user query. This shift makes inference efficiency a critical challenge for both sustainability and deployment cost.
@@
A key contributor to inference complexity in LLMs is the attention mechanism. Since the original Transformer design (Vaswani et al., 2017), the standard Multi-Head Attention (MHA) has undergone several modifications aimed at improving efficiency. Variants such as Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped-Query Attention (GQA) (Ainslie et al., 2023) reduce memory requirements by sharing key–value projections across heads, thereby lowering the cost of maintaining and accessing the Key–Value (KV) cache during decoding. Recent studies (Dao et al., 2022) further highlight that attention kernels are often memory-bound rather than compute-bound, making their design a critical factor for inference efficiency at scale. These insights underscore the central role of attention in determining both the computational and energy footprint of LLM inference.
The energy consumption of LLM inference is shaped by a complex interplay of model size, architecture, and workload geometry (i.e., the distribution of input and output sequence lengths). While recent studies have examined inference energy consumption through empirical benchmarks (Fernandez et al., 2025), they often focus on specific models (Stojkovic et al., 2024) or high-level metrics such as latency and throughput (Maliakel et al., 2025). Previous works (Wilkins et al., 2024) have assumed linear relation between the workload parameters ($n_{\text{in}}$ = input tokens, $n_{\text{out}}$ = output tokens) and the energy consumed by each generated token. Our characterization show a different result: the energy per token has local minima in $n_{\text{in}}$ and $n_{\text{out}}$ space, which can’t be explained by a linear model.
In this work, we propose an analytical model derived from the computational and memory access complexity of the Transformer, which captures the non-linear relationship between input and output sequence lengths. To evaluate our approach, we conduct a systematic analysis of LLM inference energy consumption using TensorRT-LLM, a state-of-the-art inference stack optimized for modern GPUs, running on NVIDIA H100 accelerators. Our study covers a broad set of models spanning from 1B to 9B parameters, including OPT (Zhang et al., 2022), LLaMA (Grattafiori et al., 2024), Gemma (Team et al., 2024), Falcon (Almazrouei et al., 2023), Qwen2 (Yang et al., 2024), and Granite (Mishra et al., 2024), and explores efficiency across input and output lengths ranging from 64 to 4096 tokens.
Our contributions are threefold:
@@
  • We present an analytical model that explicitly relates energy consumption to the computational cost (measured in FLOPs) of inference. This formulation reveals how quadratic input costs and linear decoding costs combine to produce the efficiency patterns observed in practice. Furthermore, we also derive an additional formulation of the model, which integrates Memory access into the equation.
@@
  • We validate these models through extensive measurements across multiple LLM families, showing that the FLOPs and Memory Access based model accurately captures the location of efficiency Sweet Spots as a function of input and output lengths, achieving a mean MAPE of 1.79% ($\pm$ 0.61).
@@
Our results suggest that production deployments can optimize Energy Efficiency by tailoring batching strategies, truncating overly long inputs, or aligning output lengths with the efficiency Sweet Spots predicted by our model.
The rest of the paper is organized as follows. After giving the necessary background in Section 2, we review related work in Section 3. Then in Sections 4 and 5 we present our analytical models and experimental methodology, before we present our experimental results in Section 6. We discuss our findings in Section 7 and conclude in Section 8.
@@ 2. Background
2.1. LLMs Architecture
LLMs are based on the Transformer architecture (Vaswani et al., 2017), which has become the standard for generative AI. A transformer block is composed of two main components: the Multi-Head self-Attention mechanism (MHA) and the Feed-Forward Network (FFN), both surrounded by residual connections, layer normalization, and dropout. The attention mechanism is commonly implemented as MHA, where the hidden representation is projected into different subspaces (heads). Each head learns to capture different types of dependencies in the sequence, and their outputs are later combined. While this improves the model’s expressivity and ability to represent diverse relationships, it also increases the total number of operations roughly in proportion to the number of heads. To reduce this overhead, Multi-Query Attention (MQA) was proposed in 2019 (Shazeer, 2019), modifying MHA by keeping multiple query heads but sharing a single set of KV projections across all heads. This design drastically reduces memory usage and KV cache size during decoding, making inference more efficient. More recently, Grouped-Query Attention (GQA) was introduced in 2023 (Ainslie et al., 2023) as a generalization of MHA and MQA: keys and values are shared within groups of heads, striking a balance between the efficiency of MQA and the expressivity of MHA.
2.2. LLMs Computation
In large-scale models, the majority of parameters reside in the FFN layers, while the attention mechanism dominates runtime complexity due to its quadratic dependence on the input sequence length. LLM inference is typically divided into two phases: prefill, where the full input sequence is processed in parallel to initialize the KV cache; and decode, where tokens are generated autoregressively. In the prefill stage, the computational cost is dominated by the quadratic attention term, whereas in the decode stage, the cost per token is linear with respect to the sequence length, due to the reuse of cached keys and values. Despite this, the decode phase often dominates total inference time for long outputs, as each token must be produced sequentially, limiting opportunities for parallelization. On modern GPUs, the throughput of LLM inference is not solely limited by arithmetic performance but also by memory bandwidth. Storing and retrieving the large KV cache during decoding creates significant pressure on the memory subsystem with high bandwidth requirements. For instance, in autoregressive decoding, each generated token requires the loading of all previous keys and values from memory, resulting in attention kernels being strongly memory-bound, particularly for long contexts (Yuan et al., 2024).
@@ 2.3. Inference Frameworks
When it comes to executing LLM inference on production hardware, several software runtimes exist, such as TensorRT-LLM (NVIDIA, 2025), vLLM (Kwon et al., 2023) and DeepSpeed (Aminabadi et al., 2022), each offering different levels of optimization and targeting different deployment constraints. These runtimes vary substantially in how they manage memory, parallelize attention, and schedule kernels. TensorRT-LLM focuses on graph-level optimizations and low-level kernel fusion, vLLM is an open-source inference system explicitly tailored for high-throughput LLM serving. Its main contribution, the PagedAttention mechanism, improves memory efficiency by partitioning key–value caches, thereby enabling scalable batch inference with reduced memory overhead. DeepSpeed, developed by Microsoft, provides a distributed training and inference framework aimed at maximizing scalability using techniques such as the Zero Redundancy Optimizer (Rajbhandari et al., 2020).
@@ 2.3.1. TensorRT-LLM
Beyond these system-level design advantages, recent empirical evidence further motivates the use of TensorRT-LLM. In particular, the authors of (Niu et al., 2025) present a comparative analysis of the frameworks mentioned above, and finds that vLLM and TensorRT-LLM consistently achieves the most energy efficient token generation among the compared frameworks, especially under high-concurrency and high-throughput workloads. Component-level measurements additionally show that both vLLM and TensorRT-LLM consumes about the $5\%$ GPU energy with respect to the consumption of a general-purpose serving frameworks as Transformers (Wolf et al., 2020), demonstrating how its hardware-aware optimizations translate directly into power savings. Collectively, these results show that TensorRT-LLM is not only engineered for performance but is also empirically validated as one of the most energy-efficient inference engines available. Based on the above consideration and to obtain production-grade figure of merit we choose TensorRT-LLM as inference run-time in this paper. Future works will extend the analysis on different ones.
@@ 2.3.2. In-Flight Batching
A key runtime optimization employed by TensorRT-LLM to maximize throughput under dynamic workloads is its in-flight batching strategy. Unlike static batching, where requests must be grouped prior to execution and processed synchronously, in-flight batching enables the runtime to continuously merge and schedule requests that are already undergoing inference. This mechanism is particularly well suited for autoregressive LLM decoding, where sequences progress token by token and naturally diverge in length. At a high level, TensorRT-LLM decouples request admission from kernel execution. Incoming inference requests are first tokenized and then admitted into an active set of sequences maintained by the runtime scheduler. During each decoding iteration, TensorRT-LLM dynamically forms micro-batches composed of all active sequences that are ready to generate the next token. Newly arriving requests can be inserted into this active set without waiting for the completion of previous batches, while completed sequences are removed as soon as they reach their termination condition. As a result, GPU execution remains continuously occupied, reducing idle cycles that would otherwise arise from batch fragmentation. From a systems perspective, in-flight batching trades strict per-request latency determinism for significantly improved throughput and hardware utilization, being especially effective in production serving scenarios, where request arrival times and sequence lengths are inherently unpredictable.
The behavior of in-flight batching is governed by user-configurable limits, notably max_batch_size and max_num_tokens. The max_batch_size parameter specifies the maximum number of active requests that can be scheduled concurrently in a decoding iteration, effectively bounding the degree of request-level parallelism. In contrast, max_num_tokens constrains the total number of tokens processed in a single micro-batch, aggregating across all active sequences. This token-level limit is particularly important for controlling GPU memory usage per iteration.
@@ 3. Related Work
Several recent works have characterized and modeled the energy footprint of LLM inference. (Fernandez et al., 2025) presents a systematic analysis of inference efficiency optimizations across diverse NLP and generative AI workloads, considering input–output token distributions, batching strategies, GPU hardware, software frameworks, and decoding methods. Their work shows that real-world energy consumption is highly sensitive to workload geometry and software–hardware configurations.
Authors of (Niu et al., 2025) characterize the power consumption of different inference engines (Transformers, vLLM, DeepSpeed, and TensorRT-LLM). They compute the Energy-per-Token metric $E_{\text{tok}}$ as:
(1) $E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}},$
where $E_{\text{tot}}$ is the total energy consumption during inference and $n_{\text{out}}$ the total generated output tokens.
In (Wilhelm et al., 2025) they investigate the trade-offs between accuracy and energy consumption when applying test-time compute strategies. They propose energy-aware routing mechanisms to guide sustainable model selection and inference, defining $E_{\text{tok}}$ as:
(2) $E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}+n_{\text{in}}}.$
Finally, in (Wilkins et al., 2024) they characterize the energy consumption and runtime behavior of a set of LLMs using A100 GPUs, proposing workload-based energy models that, for a given model, predict consumption as a function of only the input and output tokens. To the best of our knowledge, this is the only work that proposes a model to describe the total energy consumption as a function of input and output token as follows:
(3) $E_{\text{tot}}=\theta_{0}n_{\text{in}}+\theta_{1}n_{\text{out}}+\theta_{3}n_{\text{in}}n_{\text{out}}.$
In contrast to these studies, we introduce an analytical model of LLM inference energy consumption derived from both computational and memory-access complexity, enabling us to characterize the nonlinear interplay between input and output sequence lengths. To assess this model, we perform a systematic study of inference energy usage using TensorRT-LLM on NVIDIA H100 GPUs, covering a diverse set of models and spanning a wide range of sequence lengths. Our formulation makes explicit how quadratic input-processing costs and linear decoding costs combine to shape the efficiency patterns seen in practice, and we provide an extended version of the model that incorporates memory-access effects. Through extensive evaluation across multiple model families, we show that the analytical model closely reflects real energy trends and correctly identifies where efficiency peaks occur as a function of input and output lengths. Finally, we demonstrate that efficiency does not increase or decrease uniformly with sequence length but instead follows distinct regimes, with clear sweet spots and clear degradation zones depending on the balance between input size and output length.
@@
FLOPs-based (Ours) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}n_{\text{in}}+\frac{\theta_{3}n_{\text{in}}}{n_{\text{out}}}+\theta_{4}n_{\text{out}}$
Memory and FLOPs-based (Ours) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}n_{\text{in}}+\frac{\theta_{3}n_{\text{in}}}{n_{\text{out}}}+\theta_{4}n_{\text{out}}+\frac{\theta_{5}}{n_{\text{out}}}$
4. LLM Inference Consumption Analytical Models
In this section, we present our analytical frameworks for modeling LLM energy consumption, explicitly analyzing the distinct computational characteristics of the prefill and decode phases. To assess the validity of our approach, we compare our proposed models against four baselines, summarized in Table 1. From (Niu et al., 2025) we derive Baseline 1, which simply describes the $E_{\text{tok}}$ as constant. Moreover, we formalize a parametrized model that relies solely on the number of generated tokens, $n_{\text{out}}$, to characterize $E_{\text{tok}}$, referred to as Baseline 2. The third baseline, Baseline 3, expands on this by also including the number of input tokens, $n_{\text{in}}$, as introduced in Equation 2. This formulation allows us to establish a relationship between $E_{\text{tok}}$ and the total sequence length. Finally, Baseline 4 is derived directly from Equation 3.
4.1. FLOPs-Based Model
To evaluate the Energy Efficiency of our LLMs, we derived an analytical model of the computational complexity of inference. The model accounts for both the prefill phase, where the entire input sequence is processed in parallel, and the decode phase, where tokens are generated autoregressively using the cached key–value (KV) states. In this section, we present the derivation of the formulas for each phase and then combine them into a unified expression. During the prefill stage, the model processes an input sequence of length $n_{\text{in}}$ in parallel. For a single Transformer layer with hidden size $d$, the computational cost can be decomposed into two main contributions: the Self-Attention Mechanism and the Feed-Forward Network. We neglect the cost of scaling, softmax, layer normalization, dropout, and bias additions, which are negligible w.r.t. the cost of attention and FFN. Moreover, we adopt the standard FLOPs accounting convention where one multiply–accumulate operation is counted as two FLOPs (one multiplication and one addition).

### v2

# SweetSpot: An Analytical Model for Predicting Energy Efficiency of LLM Inference

@@ Abstract.
Large Language Models (LLMs) inference is central to modern AI applications, dominating worldwide datacenter workloads, making it critical to predict its energy footprint. Existing approaches estimate energy consumption as a simple linear function of input and output sequence. However, by analyzing the autoregressive structure of Transformers, which implies a fundamentally non-linear relationship between input and output sequence lengths and energy consumption, we demonstrate the existence of a generation energy minima. Peak efficiency occurs with short-to-moderate inputs and medium-length outputs, while efficiency drops sharply for long inputs or very short outputs. Consequently, we propose SweetSpot, an analytical model derived from the computational and memory-access complexity of the Transformer architecture, which accurately characterizes the efficiency curve as a function of input and output lengths. To assess accuracy, we measure energy consumption using TensorRT-LLM on NVIDIA H100 GPUs across a diverse set of LLMs ranging from 1B to 9B parameters, including OPT, LLaMA, Gemma, Falcon, Qwen2, and Granite. We test input and output lengths from 64 to 4096 tokens and achieve a mean MAPE of 1.79%. Our results show that aligning sequence lengths with these efficiency “sweet spots” reduce energy usage, up to $33.41\times$, enabling informed truncation, summarization, and adaptive generation strategies in production systems.
LLM Energy Consumption; Energy Efficiency Modeling; LLM Sustainability
@@ 1. Introduction
Large Language Models (LLMs) have rapidly emerged as the foundation of modern Natural Language Processing (NLP), enabling applications ranging from reasoning and summarization to dialogue systems and generative Artificial Intelligence (AI). Their adoption at scale, however, has introduced a growing concern: the substantial energy consumption associated with inference, which has been estimated to account for up to 90% of the model’s life-cycle energy consumption (Jegham et al., 2025). Unlike training, which is an episodic process, inference occurs continuously in production settings, every time an LLM generates output for a user query. This shift makes inference efficiency a critical challenge for both sustainability and deployment cost.
@@
A key contributor to inference complexity in LLMs is the attention mechanism. Since the original Transformer design (Vaswani et al., 2017), the standard Multi-Head Attention (MHA) has undergone several modifications aimed at improving efficiency. Variants such as Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped-Query Attention (GQA) (Ainslie et al., 2023) reduce memory requirements by sharing Key–Value (KV) projections across heads, thereby lowering the cost of maintaining and accessing the KV cache during decoding. Recent studies (Dao et al., 2022) further highlight that attention kernels are often memory-bound rather than compute-bound, making their design a critical factor for inference efficiency at scale. These insights underscore the central role of attention in determining both the computational and energy footprint of LLM inference.
The energy consumption of LLM inference is shaped by a complex interplay of model size, architecture, and workload geometry (i.e., the distribution of input and output sequence lengths), and its characterization requires thorough empirical validation.(Fernandez et al., 2025) Recent studies on inference energy consumption focus only on specific models (Stojkovic et al., 2024) or coarse-grained metrics (e.g., latency and throughput (Maliakel et al., 2025)), which fail to generalize and provide fine-grained inference energy profiling. Moreover, previous claims (Wilkins et al., 2024) on linear relation between the workload parameters ($n_{\text{in}}$ = input tokens, $n_{\text{out}}$ = output tokens) leads to the conclusion that every generated token incurs a constant computational and energy cost, as reflected by commercial AI providers that charge inference services based on the number of generated output tokens or a linear combination of input and output tokens. However, transformer structure and autoregressive generation process, with distinct prefill and decode phases, suggest a more complex dependency of computational cost, and consequently energy consumption, on both input and output sequence lengths. This complexity is not fully captured by current state-of-the-art analytical models.
To address these limitations, we conduct a thorough and systematic analysis of the energy consumption of a broad set of LLMs (from 1B to 9B parameters), and we explore efficiency across a wide range of input and output lengths (from 64 to 4096 tokens). Following the results of our investigation, in this paper we present the following contributions:
@@
  • The formulation of SweetSpot, an analytical model derived from computational and memory-access complexity of inference, capable of predicting the peak energy efficiency input and output lengths configuration. The proposed model reveals how quadratic input costs and linear decoding costs combine to produce the efficiency patterns observed in practice.
@@
  • We validate SweetSpot through extensive measurements across multiple LLM families, showing that it accurately captures the location of peak energy efficiency configurations (“sweet spots”) as a function of input and output lengths, achieving a mean MAPE of 1.79% ($\pm$ 0.61).
@@
We conduct our study using TensorRT-LLM, [^1] https://github.com/NVIDIA/TensorRT-LLM a state-of-the-art inference stack optimized for modern GPUs, running on NVIDIA H100 accelerators. Through TensorRT-LLM, we test several widely-used LLM models, such as OPT (Zhang et al., 2022), LLaMA (Grattafiori et al., 2024), Gemma (Team et al., 2024), Falcon (Almazrouei et al., 2023), Qwen2 (Yang et al., 2024), and Granite (Mishra et al., 2024). Doing so allows us to obtain reliable and realistic results which apply to a broad range of different models and architectures.
Our results suggest that production deployments can optimize energy efficiency by tailoring batching strategies, truncating overly long inputs, or aligning output lengths with the peak energy efficiency configurations predicted by our model.
The rest of the paper is organized as follows. After giving the necessary background in Section 2, we review related work in Section 3. Then, in Section 4, we present our analytical model SweetSpot. In Section 5 we explain our experimental methodology and illustrate the experimental results. Finally, we discuss our current limitations and future works in Section 6, and conclude in Section 7.
@@ 2. Background
This section provides the necessary background for the analysis presented in this work. We briefly review the fundamentals of LLM architectures, their computational characteristics, and the main design principles of modern inference frameworks.
2.1. LLM Architecture
LLMs are based on the Transformer architecture (Vaswani et al., 2017), which has become the standard for generative AI. A transformer block is composed of two main components: the Multi-Head self-Attention mechanism (MHA) and the Feed-Forward Network (FFN), both surrounded by residual connections, layer normalization, and dropout. The attention mechanism is commonly implemented as MHA, where the hidden representation is projected into different subspaces (heads). Each head learns to capture different types of dependencies in the sequence, and their outputs are later combined. While this improves the model’s expressivity and ability to represent diverse relationships, it also increases the total number of operations roughly in proportion to the number of heads. To reduce this overhead, Multi-Query Attention (MQA) was proposed in 2019 (Shazeer, 2019), modifying MHA by keeping multiple query heads but sharing a single set of Key-Value (KV) projections across all heads. This design drastically reduces memory usage and KV cache size during decoding, making inference more efficient. More recently, Grouped-Query Attention (GQA) was introduced in 2023 (Ainslie et al., 2023) as a generalization of MHA and MQA: keys and values are shared within groups of heads, striking a balance between the efficiency of MQA and the expressivity of MHA.
2.2. LLM Inference Computation
In large-scale models, the majority of parameters reside in the FFN layers, while the attention mechanism dominates runtime complexity due to its quadratic dependence on the input sequence length.
LLM inference is typically divided into two phases: prefill, where the full input sequence is processed in parallel to initialize the KV cache; and decode, where tokens are generated autoregressively (Yuan et al., 2024). In the prefill stage, the computational cost is dominated by the quadratic attention term, whereas in the decode stage, the cost per token is linear with respect to the sequence length, due to the reuse of cached keys and values. Despite this, the decode phase often dominates total inference time for long outputs, as each token must be produced sequentially, limiting opportunities for parallelization.
On modern GPUs, the throughput of LLM inference is not solely limited by arithmetic performance but also by memory bandwidth. Storing and retrieving the large KV cache during decoding creates significant pressure on the memory subsystem with high bandwidth requirements. For instance, in autoregressive decoding, each generated token requires the loading of all previous keys and values from memory, resulting in attention kernels being strongly memory-bound, particularly for long contexts.
@@ 2.3. Inference Frameworks
When executing LLM inference on production hardware, several software for inference run-time exist, such as TensorRT-LLM (NVIDIA, 2025), vLLM (Kwon et al., 2023) and DeepSpeed (Aminabadi et al., 2022), each offering different levels of optimization and targeting different deployment constraints. They vary substantially in how they manage memory, parallelize attention, and schedule kernels. TensorRT-LLM focuses on graph-level optimizations and low-level kernel fusion, while vLLM is an open-source inference system explicitly tailored for high-throughput LLM serving. Its main contribution, the PagedAttention mechanism, improves memory efficiency by partitioning KV caches, thereby enabling scalable batch inference with reduced memory overhead. DeepSpeed, developed by Microsoft, provides a distributed training and inference framework aimed at maximizing scalability using techniques such as the Zero Redundancy Optimizer (Rajbhandari et al., 2020).
@@ 2.3.1. TensorRT-LLM
Beyond these system-level design advantages, recent empirical evidence further motivates the use of TensorRT-LLM. In particular, the authors of (Niu et al., 2025) present a comparative analysis of the frameworks mentioned above, and find that vLLM and TensorRT-LLM consistently achieve the most energy-efficient token generation among the compared frameworks, especially under high-concurrency and high-throughput workloads. Component-level measurements further show that both vLLM and TensorRT-LLM consume only about $5\%$ of the GPU energy compared to a general-purpose serving framework such as Transformers (Wolf et al., 2020). This result highlights how their hardware-aware optimizations translate directly into power savings. Collectively, these results show that TensorRT-LLM is not only engineered for performance but is also empirically validated as one of the most energy-efficient inference engines available. Based on these considerations, as well as to obtain a production-grade figure of merit, we choose TensorRT-LLM as the inference run-time in this paper. Future work will extend the analysis to different frameworks.
@@ 2.3.2. In-Flight Batching
A key run-time optimization employed by TensorRT-LLM to maximize throughput under dynamic workloads is its in-flight batching strategy. Unlike static batching, where requests must be grouped prior to execution and processed synchronously, in-flight batching enables the run-time to continuously merge and schedule requests that are already undergoing inference. This mechanism is particularly well-suited for autoregressive LLM decoding, where sequences progress token by token and naturally diverge in length. At a high level, TensorRT-LLM decouples request admission from kernel execution. Incoming inference requests are first tokenized and then admitted into an active set of sequences maintained by the run-time scheduler. During each decoding iteration, TensorRT-LLM dynamically forms micro-batches composed of all active sequences that are ready to generate the next token. Newly arriving requests can be inserted into this active set without waiting for the completion of previous batches, while completed sequences are removed as soon as they reach their termination condition. As a result, GPU execution remains continuously occupied, reducing idle cycles that would otherwise arise from batch fragmentation. From a system perspective, in-flight batching trades strict per-request latency determinism for significantly improved throughput and hardware utilization, being especially effective in production serving scenarios, where request arrival times and sequence lengths are inherently unpredictable.
The behavior of in-flight batching is managed by user-configurable limits, namely max_batch_size and max_num_tokens. The max_batch_size parameter specifies the maximum number of active requests that can be scheduled concurrently in a decoding iteration, effectively bounding the degree of request-level parallelism. In contrast, max_num_tokens constrains the total number of tokens processed in a single micro-batch, aggregating across all active sequences. This token-level limit is particularly important for controlling GPU memory usage per iteration.
@@ 3. Related Work
Several recent works have characterized and modeled the energy footprint of LLM inference. For instance, (Fernandez et al., 2025) presents a systematic analysis of inference energy-efficiency optimizations across diverse NLP and generative AI workloads, considering input–output token distributions, batching strategies, GPU hardware, software frameworks, and decoding methods. Their work shows that real-world energy consumption is highly sensitive to workload geometry and software–hardware configurations. Authors of (Niu et al., 2025) characterize the energy consumption of different inference engines (Transformers, vLLM, DeepSpeed, and TensorRT-LLM). They compute the Energy-per-Token metric $E_{\text{tok}}$ as:
(1) $E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}}$
where $E_{\text{tot}}$ is the total energy consumption during inference and $n_{\text{out}}$ the total generated output tokens. In (Wilhelm et al., 2025), they investigate the trade-offs between accuracy and energy consumption when applying test-time compute strategies. They propose an energy-aware routing mechanisms to guide sustainable model selection and inference, defining $E_{\text{tok}}$ to account for the number of input tokens $n_{\text{in}}$, by expressing it as:
(2) $E_{\text{tok}}=\frac{E_{\text{tot}}}{n_{\text{out}}+n_{\text{in}}}$
Finally, in (Wilkins et al., 2024) they characterize the energy consumption and run-time behavior of a set of LLMs using A100 GPUs, proposing workload-based energy models that, a given model predicts consumption as a function of only the input and output tokens. To the best of our knowledge, this is the only work that proposes a model to describe the total energy consumption as a function of input and output token as follows:
(3) $E_{\text{tot}}=\theta_{0}n_{\text{in}}+\theta_{1}n_{\text{out}}+\theta_{3}n_{\text{in}}n_{\text{out}}$
In contrast to prior studies, we introduce SweetSpot, an analytical model of LLM inference energy consumption derived from both computational and memory-access complexity. This model enables us to characterize the non-linear interplay between input and output sequence lengths. To assess it, we conduct a systematic study of inference energy usage using TensorRT-LLM on NVIDIA H100 GPUs, across a diverse set of models and a wide range of sequence lengths. Our formulation makes explicit how quadratic input-processing costs and linear decoding costs combine to shape the efficiency patterns observed in practice, and we further extend the model to incorporate memory-access effects. Through extensive evaluation across multiple model families, we show that SweetSpot closely matches measured energy trends and correctly identifies where efficiency peaks occur as a function of input and output lengths. Finally, we show that efficiency does not vary monotonically with sequence length; instead, it follows distinct regimes, with clear ”sweet-spots” configurations and degradation zones depending on the balance between input size and output length.
@@
SweetSpot (FLOPs-only) (ours) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}n_{\text{in}}+\frac{\theta_{3}n_{\text{in}}}{n_{\text{out}}}+\theta_{4}n_{\text{out}}$
SweetSpot (ours) $E_{\text{tok}}(n_{\text{in}},n_{\text{out}})=\theta_{0}+\frac{\theta_{1}n_{\text{in}}^{2}}{n_{\text{out}}}+\theta_{2}n_{\text{in}}+\frac{\theta_{3}n_{\text{in}}}{n_{\text{out}}}+\theta_{4}n_{\text{out}}+\frac{\theta_{5}}{n_{\text{out}}}$
4. SweetSpot: Model Formulation and Energy Efficiency Prediction
In this section, we present SweetSpot, our analytical model for predicting LLM inference energy efficiency, explicitly analyzing the distinct computational characteristics of the prefill and decode phases. We first introduce a version of SweetSpot based solely on the computational complexity of inference, and subsequently extend it to incorporate memory-access complexity. To assess the validity of our approach, we compare SweetSpot against four baselines, summarized in Table 1. From (Niu et al., 2025) we derive Baseline 1, which simply describes the $E_{\text{tok}}$ as constant. Moreover, we formalize a parametrized model that relies solely on the number of generated tokens, $n_{\text{out}}$, to characterize $E_{\text{tok}}$, referred to as Baseline 2. The third baseline, Baseline 3, expands on this by also including the number of input tokens, $n_{\text{in}}$, as introduced in Equation 2. This formulation allows us to establish a relationship between $E_{\text{tok}}$ and the total sequence length. Finally, Baseline 4 is derived from Equation 3, normalizing the equation by $n_{\text{out}}$.
4.1. SweetSpot Model (FLOPs-only)
To evaluate the energy efficiency of our LLMs, we derived an initial analytical model grounded exclusively in the computational complexity of inference. The model accounts for both the prefill phase, where the entire input sequence is processed in parallel, and the decode phase, where tokens are generated autoregressively using the cached KV states. In the following, we present the derivation of the formulas for each phase and then combine them into a unified expression. During the prefill stage, the model processes an input sequence of length $n_{\text{in}}$ in parallel. For a single Transformer layer with hidden size $d$, the computational cost can be decomposed into two main contributions: the Self-Attention Mechanism and the Feed-Forward Network (FFN). We neglect the cost of scaling, softmax, layer normalization, dropout, and bias additions, which are negligible w.r.t. the cost of attention and FFN.
Moreover, we adopt the standard Floating-point Operations (FLOPs), denoted by $F$, accounting convention where one multiply–accumulate operation is counted as two FLOPs (one multiplication and one addition).